OpenStudy (anonymous):

A=1/2h(B+b) .... when A=60, B=7, and b=3

OpenStudy (anonymous):

60=1/2h(7+3) 60=1/2h(10) divide by 10 1/2h=6 divide by 1/2/multiply by 2 h=12